OpenStudy (anonymous):

Two birds sit at the top of two different trees 17.8 feet away from one another. The distance between the first bird and a bird watcher on the ground is 37.4 feet. What is the angle measure, or angle of depression, between this bird and the bird watcher? A. 25.5° B. 28.4° C. 61.6° D. 64.5°

OpenStudy (anonymous):

This question is weird. what does the 2nd bird has anything to do at all. since the question only involves the first bird and the bird watcher.

OpenStudy (anonymous):

No wonder. without the diagram this would be unsolvable. cos x = adj/hyp cos x = 17.8/37.4= 0.475935829 then just cos^-1 (0.476) = 61.6 degrees approx.